Two sides of a triangle are equal in length. The length of the third side is three meters less than the sum of the lengths of the two other sides. Find the length of the longest side of the triangle if the perimeter is 29 meters.
Answer by mananth(16946)

Two sides of a triangle are equal in length. The length of the third side is three meters less than the sum of the lengths of the two other sides. Find the length of the longest side of the triangle if the perimeter is 29 meters.
...
one side = x
other side =x
third side= 2x-3
..
x+x+2x-3=29
4x-3=29
+3 +3
4x=32
/4
x=8
x=8 meters
Longest side = 2x-3 = 16-3=13 meters
...
CHECK
8+8+13=29 meters
...
m.ananth@hotmail.ca